OpenStudy (anonymous):

Can someone help me find the derivative of the problem below? I'll give a medal! :)

OpenStudy (anonymous):

\[\frac{ 5 }{ (2x)^3 }\]

OpenStudy (sleepyhead314):

do you want to use quotient rule or product rule? personally I think product rule is easier xD

OpenStudy (anonymous):

Whatever you think is easier, I'm okay with.

OpenStudy (sleepyhead314):

mmmkk well do you know product rule? :) \[\frac{ d }{ dx } [f(x) * g(x)] ~~=~~ f(x)*g'(x) + g(x)*f'(x)\]right? :)

OpenStudy (anonymous):

Yes, I am familiar with it

OpenStudy (sleepyhead314):

mmm I'm not awake we can rewrite the given equation as \[5*(2x)^{-3}\]right?

OpenStudy (anonymous):

Yes

OpenStudy (sleepyhead314):

turns out we don't need product rule xD (cuz the numeator is a constant...) ingore what me said before we use power rule and chain rule \[5*(2x)^{-3} \rightarrow 5(-3)*(2x)^{-4}*(\frac{ d }{ dx }2x)\]

OpenStudy (sleepyhead314):

tell me if you get confused xD

OpenStudy (anonymous):

Sorry, my page keeps reloading. And that makes sense

OpenStudy (anonymous):

You'd get -30(-2x)^4

OpenStudy (anonymous):

-30(2x)^4 sorry the 2s not negative

OpenStudy (sleepyhead314):

mhmm then we simplify and get \[-15 * (2x)^{-4} * (2) \rightarrow -30 * (2x)^{-4} \rightarrow \frac{ -30 }{ (2x)^{4} } \]\(\huge \rightarrow \frac{-30}{16x^4} \rightarrow \frac{-15}{8x^4}\)

OpenStudy (freckles):

you can also just use power rule... if you write \[5(2x)^{-3} =5(2)^{-3}x^{-3} =\frac{5}{2^3}x^{-3}=\frac{5}{8}x^{-3} \\ \text{ then differentiate using power rule }\]

OpenStudy (sleepyhead314):

yay math thanks @freckles :)

OpenStudy (anonymous):

Thank you, both of you. :)

OpenStudy (sleepyhead314):

glad we could helppp :)